ESTIMATION OF THE SIZE AND MEAN VALUE OF A STIGMATIZED CHARACTERISTIC OF A HIDDEN GANG IN A FINITE POPULATION:

A UNIFIED APPROACH

RAGHUNATH ARNAB^{1} AND SARJINDER SINGH^{2}

1Department of Statistics, University of Durban-Westville, Private Bag-X54001, Durban-4000,

South Africa, e-mail: arnab@pixie.udw.ac.za

2Department of Mathematics and Statistics, University of Saskatchewan, Saskatoon, SK, Canada S7N 5E6, e-mail: sarjinder@yahoo.com

(Received May 1, 2000; revised April 25, 2001)

Abstract. We consider the problem of estimating the size and the mean value of a stigmatized quantitative characteristic of a sub-group (or hidden gang) in a finite population using a unified approach. The proposed method may be useful in estimating the proportion and the mean income of terrorists, “hijackers” or “freedom fighters” of a particular country, including those who operate across different countries.

Key words and phrases: Randomized response, sampling strategies, estimation of proportion, estimation of mean.